Hydrogen gas can be collected from the reaction of a metal element with a strong acid as

shown. When a 500.0 mg iron metal sample was soaked in 50.00 mL 0.500 M HNO3, 16.00 mg H2 was collected. How efficient is this process?

2Fe + 6HNO3 → 2Fe(NO3)3 + 3H2

1
Expert's answer
2021-11-13T01:33:38-0500

From the question, we have 500mg of iron which is (500/1000)g which is 0.5g

And for HNO3 we have to get the mass

So we first get the no of moles which is;

n=CV

n=0.5×0.05

n=0.025

Mass=mole×molar mass

Mass=0.025×63

Mass=1.575g

So now we'll find the limiting reagent

If 112g of Iron react with 378g of HNO3

0.5g- - - - - - - - - - - - - - - Xg of HNO3

X=(378×0.5)/112

X=1.69g

Therefore HNO3 is the limiting reagent

If 378g of HNO3 produces 6g of H2

1.575g - - - - - - - - - - - - - - Xg of

X=0.025g

Efficiency =(actual yield/theory yield)×100

=(0.016/0.025)×100

Efficiency=64%